• Main INDEX
  • Monthly INDEX
  • PREV
  • NEXT

    User name X. Zheng

    Log entry time 10:03:52 on July16,2003

    Entry number 104034

    keyword=Target temperature test

    Since it will take a while to bring back the Septum, we decided to do target temperature test.

    8:17am, target masked and moved to pick-up coil position;
    8:20am, NMR with three longitudinal lasers on, RF 3V rms; (forgot to print out the NMR settings); analysis gives 40.85
    8:22am, turn off lasers 4,6,7;
    8:44am, NMR with lasers off; analysis gives 41.37%;
    8:47am, turn on laser 4;
    8:53am, turn on laser 7;
    8:57am, turn on laser 6;
    9:17am, NMR with three lasers on, now RTD6 gives 210 C (instead of 214C before); analysis gives 40.38%;
    9:18am, turn off three lasers;
    9:38am, NMR with lasers off, analysis gives 40.71%;
    9:40-9:43am, turn on laser 4, 6, 7;
    10:03am, NMR with lasers on, analysis gives 39.83%;
    end